6^(3x-1)=7^(x+1)
I have tried logging both sides and then moving the exponents in front I think I'm supposed to distribute, but I can't quite remember how.
so this is where I am at:
3x-1 log 6 = x+1 log 7

Answer by ankor@dixie-net.com(22740)

3x-1 log 6 = x+1 log 7
Find the log of 6 and the log of 7, and you have
.778(3x-1) = .845(x+1)
2.334x - .778 = .845x + .845
2.334x - .845x = .845 + .778
let you do the math here
:
Check your solution in the original equation